# Non-Atomic Arbitrage in Decentralized Finance

### 15 Jan 2024 v2

@@ Abstract
In our study of non-atomic arbitrage, we uncover that more than a fourth of the volume on Ethereum’s biggest five DEXes from the merge until 31 October 2023 can likely be attributed to this type of MEV. We further highlight that only eleven searchers are responsible for more than 80% of the identified non-atomic arbitrage volume sitting at a staggering 132 billion US\$ and draw a connection between the centralization of the block construction market and non-atomic arbitrage. Finally, we discuss the security implications of these high-value transactions that account for more than 10% of Ethereum’s total block value and outline possible mitigations.
@@ 1 Introduction
The introduction of decentralized exchanges (DEXes) has made the trading of cryptocurrencies a centerpiece of decentralized finance (DeFi). Today, a multitude of DEXes exist on several different chains. The prices on each of these DEXes are the product of previous transactions on the respective DEX and thus can differ from the prices quoted at a given time on centralized exchanges (CEX) or other DEXes. These differences represent an arbitrage opportunity, which a trader can exploit by buying the asset at the cheaper venue and selling it at the other. We refer to price differences between DEXes on the same chain as atomic arbitrage, since these arbitrages can be carried out within the same transaction, whereas non-atomic arbitrage stems from price discrepancies involving DEXes on two different chains or, presumably more common, centralized exchanges (CEX). The prior is known as cyclic arbitrage and has been extensively studied [1]. We focus on the latter, which is naturally more challenging to analyze as it involves actions outside the respective blockchain. We note that while non-atomic arbitrage opportunities exist on various chains we focus on Ethereum, as it is home to the by far largest portion of DeFi applications [2].
@@
In this work, we present the first in-depth analysis of non-atomic arbitrage on Ethereum DEXes to show that non-atomic arbitrage, a type of MEV, accounts for a significant proportion of the total DEX volume. We further identify alarming centralizing trends in the block-building market as a direct consequence of non-atomic arbitrage.
@@ 2.3 Decentralized Exchanges
DEXes are one of the most popular DeFi applications. They allow users to exchange cryptocurrencies on the blockchain without giving up custody of their assets. While traditional CEXes generally utilize a limit order book to match orders, most DEXes are constant function market makers (CFMM). With a CFMM, users execute their order against a liquidity pool holding reserves of two or more cryptocurrencies. Trades against the pool execute according to a pre-defined trading function specified in the pool’s smart contract, i.e., the pool only accepts trades that keep its trading function constant. Importantly, every trade that executes in a CFMM pool moves the pool’s price deterministically, dependent only on the pool’s reserves ahead of the trade and the trade’s input size. Consequently, if an arbitrage opportunity arises, the assets locked in the liquidity pool determine how large the arbitrage trade can be. Hence, the larger the liquidity pools, the larger the required arbitrage trades are to synchronize the prices across different trading venues.
@@ 3.1 Ethereum Blockchain
Maximal Extractable Value. To differentiate HFT trades from types of MEV and to compare their respective prevalence on DEXes, we collect MEV transaction labels from zeromev [9]. These labels specify whether a transaction was part of a sandwich attack (i.e., victim and attacker transaction), performs a cyclic arbitrage, or a liquidation.
@@ 3.2 PBS Relays
We collect data from eleven relays: Aestus, Agnostic, Blocknative, bloXroute (Ethical), bloXroute (Max Profit), bloXroute (Regulated), Eden, Flashbots, Manifold, Relayoor and UltraSound. The relays implement public APIs that give access to the blocks delivered by the relays to the proposers, as well as the bids received by the builders. We note that some relay APIs, namely Blocknative, bloXroute (Ethical), Eden, and Relayoor, are no longer reachable. For these, our data only extends to 31 April 2023, i.e., the last time we scraped the data. Further, bloXroute (Ethical), bloXroute (Max Profit), and bloXroute (Regulated) removed access for builder bids from their APIs. When querying the builder bids API for blocks that passed through the respective bloXroute relay, the response is consistently empty. Thus, for the three bloXroute relays the bid data only extends to 31 January 2023, i.e., the last time we scraped that data and received non-empty responses.
@@ 4 Non-Atomic Arbitrage on DEXes Model
We illustrate the process a non-atomic arbitrageur follows to get the on-chain leg of the trade included in Figure 2. Upon detecting an arbitrage opportunity, the searcher submits the desired transaction to the builder. The builder then builds their block and sends it together with their bid to the relay. Next, if requested, the relay sends the highest value block to the proposer who in turn picks the highest bidding block to propose to the network. Importantly, even though visualized separately one entity could for instance operate a builder and a searcher in this case we would refer to the latter as an integrated seacher.
@@ 4.1 Non-Atomic Arbitrageur Model
We develop a simple model for the profit of a non-atomic arbitrageur. We focus on two currencies $X$ and $Y$ whose price on a DEX on Ethereum is $P_{\mathrm{on}}$ and the off-chain price $P_{\mathrm{off}}$. As discussed in Section 2.3, most DEXes are a CFMM with the largest DEX being Uniswap [8]. Uniswap currently has two active versions, V2 [14] and V3 [15]. Here, we focus on the V2 version, which is also employed by the DEX Sushiswap [16], but the model also holds for V3 between two ticks.
The CFMM for Uniswap imposes, $x\cdot y=L^{2}$, where $x$ and $y$ stand for the number of $X$- and $Y$-tokens locked in the liquidity pool, and $L$ is referred to as the liquidity. The (marginal) price $P$ for $X$-tokens in terms of $Y$-tokens is given as $P=y/x$. The reserves can be expressed in terms of the liquidity and price by solving for $x$ and $y$, yielding $x=L/\sqrt{P}$, $y=L\sqrt{P}$.
We denote the price difference between the off-chain and on-chain as $\Delta P=P_{\mathrm{off}}-P_{\mathrm{on}}$. Without loss of generality, we assume the searcher discovers $\Delta P>0$, i.e., the price of $X$-tokens off-chain is higher such that the arbitrageur buys $X$-tokens on the DEX. During the trade, they pay $\Delta y$ of which $f\cdot\Delta y$ is paid in fees to the liquidity providers, where $f$ is the protocol fee. With this and the constant product $L^{2}=xy$, the amount of $X$-tokens the trader receives is given by
@@
As long as the price off-chain remains higher than the on-chain price received by the arbitrageurs, they can extract more value by buying more $\Delta x$ on the DEX. Thus, the arbitrageurs will trade the price on the DEX until the price they receive for the last unit of $X$ is $P_{\mathrm{off}}$. The pool’s marginal price at the end of the trade is then given by
$P_{\mathrm{end}}=P_{\mathrm{off}}\cdot(1-f)=\tfrac{y+(1-f)\Delta y}{x-\Delta x}.$
Figure 3: Difference between the amount the arbitrageur pays on-chain and what is received off-chain as a function of the price difference. We set $L=10^{6}$, $P_{\mathrm{on}}=1$ and $f=0.3\%$.
Simultaneously, to this buy order, the arbitrageur sells the same amount of $X$-tokens, $\Delta x$ through the off-chain exchanges at price $P_{\mathrm{off}}$ [^1] We assume the price off-chain is not impacted by this arbitrage. . Thus, the arbitrageurs receives
$\Delta y_{\mathrm{off}}=P_{\mathrm{off}}\cdot\Delta x=\tfrac{y+(1-f)\Delta y}{x-\Delta x}\tfrac{1}{1-f}\cdot\tfrac{x(1-f)\Delta y}{y+(1-f)\Delta y},$
@@
$\Delta y_{\mathrm{off}}-\Delta y=L\tfrac{(\sqrt{(1-f)(\Delta P+P_{\mathrm{on}})}-\sqrt{P_{\mathrm{on}}})^{2}}{(1-f)\sqrt{P_{\mathrm{on}}}},$
where we also expressed $x$ and $y$ in terms of $L$ and the starting on-chain price $P_{\mathrm{on}}$. Here, we ignored potential fees payable on the off-chain exchange as their fees for large traders is typically significantly lower than the Uniswap fee. The profit as a function of the price change $\Delta P$ is plotted in Figure 3. The profit becomes significant for larger price changes, showing that sufficient price difference will attract non-atomic arbitrage trades. Note that profitable arbitrage only exists for $P_{\mathrm{on}}<P_{\mathrm{off}}\cdot(1-f)$, as otherwise the fees paid to the liquidity pool exceed the price difference that can be profited from.
While an arbitrageur still needs to pay the builder to have their trade included, the result shows how valuable the arbitrage trade is for the searcher and limits the bid they would be willing to submit.
@@ 4.2 Case study of Block 18,360,789
To provide a better understanding of non-atomic arbitrage, we go through a case study of block 18,360,789 – a block with a significant price change in the lead-up to the block, i.e., the time between the previous block proposal and the block proposal itself. In Figure 4, we plot the time of the and the value of bids from the builders. Note that we draw out bids from rsyncbuilder (shown in red) and beaverbuild (shown in blue), as these two builders were previously identified as having integrated searchers that perform these non-atomic arbitrage trades [4]. The bids from all remaining builders are shown in yellow. We further indicate the time of the bid corresponding to the block that was chosen by the proposer, i.e., the relays deliver the highest bid to the proposer when the proposer requests it. Finally, we also plot the relative price change of ETH-USDT and BTC-USDT on Binance.com during the same time. Recall that USDT is a stablecoin pegged to the US\$.
One immediately notices that the time in the lead-up to the block is characterized by extremely high price volatility of both ETH and BTC in comparison to USDT. The relative price change exceeds 0.6% and 1.3% in less than ten seconds respectively. This price change creates an arbitrage opportunity, as the prices on DEXes do not move in-between blocks. Thus, as outlined in the previous section, the first transaction in a DEX pool can close this price difference and thereby profit from the arbitrage opportunity.
@@
tx index searcher pool tokenIn tokenOut amount [US\$] 0 rsyncsearcher3 0x88e6a...f5640 USDC ETH 2148195.308 1 rsyncsearcher3 0x99ac8...abc35 USDC BTC 659507.701 2 rsyncsearcher3 0x4585f...a20c0 ETH BTC 554524.411 3 rsyncsearcher3 0x9db9e...8425b USDT BTC 186858.044 4 rsyncsearcher3 0x11b81...697f6 USDT ETH 523225.597 5 rsyncsearcher3 0x4e68c...dfa36 USDT ETH 524150.487 6 rsyncsearcher3 0xcbcdf...d62ed ETH BTC 495617.185 8 searcher1 0x8ad59...6e6d8 USDC ETH 562787.652 9 rsyncsearcher3 0x60594...5a270 DAI ETH 116090.952 10 rsyncsearcher3 0xc2e9f...a25f8 DAI ETH 105871.055 12 searcher1 0xd51a4...aae46 USDT BTC 17463.007 14 rsyncsearcher3 0x391e8...86649 DAI BTC 9214.814 17 rsyncsearcher3 0x06da0...84553 USDT ETH 18283.993 18 rsyncsearcher3 0xc3d03...5882f DAI ETH 10533.977 19 rsyncsearcher3 0x397ff...7aca0 USDC ETH 14457.825 21 rsyncsearcher3 0xceff5...d3a58 ETH BTC 15450.829 22 rsyncsearcher3 0x824a3...c0673 ETH BLUR 11451.829 23 rsyncsearcher3 0x56534...a83b2 USDT BTC 4822.975 27 rsyncsearcher3 0xcbfb0...7ef3c USDC BTC 1538.249 28 rsyncsearcher3 0x290a6...f5b42 ETH MATIC 14629.722 31 searcher5 0xcd828...a9fbf FXS ETH 6247.832 33 rsyncsearcher3 0x919fa...daf79 ETH CRV 4954.943 35 rsyncsearcher3 0x34704...08d60 USDT UNI 1106.016 36 rsyncsearcher3 0xa3f55...dfd74 LDO ETH 9727.578 43 rsyncsearcher3 0x9febc...c82ae USDC 1INCH 516.780 44 rsyncsearcher3 0x1d420...bd801 ETH UNI 6336.821
@@
TABLE I: Non-atomic arbitrages identified in block 18,360,789 on 16 October 2023. There are 26 non-atomic arbitrage swaps with a total volume of 6,023,565 US\$ executed by merely three searchers (cf. Appendix D for a mapping from address to searcher name). Further, rsyncbuilder’s integrated searcher is responsible for 23 of these arbitrages. We indicate the tx index, the searcher, pool, and the tokenIn (the token that was sold), the tokenOut (the token that was bought), and the amount (the trade volume in US\$).
@@
Coming back to Figure 4, we can observe that around five seconds after the price starts to change on Binance.com the bids start to increase. At this point, the price difference appears to be big enough for non-atomic arbitrage to be profitable. Further, we find that bids from the builders that are associated with non-atomic arbitrage transactions are higher than those from the rest and that the bids continue to increase as the ETH and BTC prices on Binance.com increase. The block chosen by the proposer was built by rsyncbuilder and its bid, i.e., the value received by the proposer was a staggering 10.32 ETH ($\approx$ 23,000 US\$ at the time of this writing). In Table I, we take an in-depth look a the non-atomic arbitrage trades we identified with our heuristics that we will introduce in the following Section 5. In the block, there were 26 transactions with volume exceeding 6 million US\$ that we identified as performing non-atomic arbitrage trades. Remarkably, all but three of these transactions were by the rsyncsearcher3 – a searcher we identified to be linked to the rsyncbuilder that won the block (cf. Section 6.3). Additionally, we highlight that the rsyncsearchers paid a remarkable 11.15 ETH (i.e., more than the PBS bid) for their transactions. Thus, the rsyncbuilder built a high-value block largely attributed to fees paid by trades we labeled as non-atomic arbitrage by its own integrated searcher.
@@
Note that most of trades buy ETH or BTC for a stablecoin. As we saw in Figure 4 the prices of these two cryptocurrencies rose on Binance.com, and they were thus trading for cheaper on DEXes at the beginning of the block. For example, if we look at the price at the beginning of the block, we see that the price of the ETH-USDT Uniswap V3 pool was 1563.57, and at the end of the block, it was 1574.15. The Binance.com price was 1564.61 at the beginning of the block’s slot, while it rose to 1574.63 by the time the block was proposed. Notice that the arbitrageurs drove the price almost exactly to the off-chain price, as we show to be optimal in our previous analysis (cf. Section 4.1). Additionally, some transactions exchange cryptocurrencies for cryptocurrencies (i.e., no stablecoins), such as the transaction at index 2 and 6, which exchange ETH for BTC. Note that the relative price increase measured in USDT of BTC was higher than that of ETH. Thus, BTC relative to ETH was also available for cheaper on DEXes at the beginning of the block. Hence, the searcher also swaps ETH for BTC.

### 08 Apr 2024 v3

@@ Abstract
In our study of non-atomic arbitrage, we uncover that more than a fourth of the volume on Ethereum’s biggest five DEXes from the merge until 31 October 2023 can likely be attributed to this type of MEV. We further highlight that only eleven searchers are responsible for more than 80% of the identified non-atomic arbitrage volume sitting at a staggering \$132 billion and draw a connection between the centralization of the block construction market and non-atomic arbitrage. Finally, we discuss the security implications of these high-value transactions that account for more than 10% of Ethereum’s total block value and outline possible mitigations.
@@ 1 Introduction
The introduction of decentralized exchanges (DEXes) has made the trading of cryptocurrencies a centerpiece of decentralized finance (DeFi). Today, a multitude of DEXes exist on several different chains. The prices on each of these DEXes are the product of previous transactions on the respective DEX and thus can differ from the prices quoted at a given time on centralized exchanges (CEX) or other DEXes. These differences represent an arbitrage opportunity, which a trader can exploit by buying the asset at the cheaper venue and selling it at the other. We refer to price differences between DEXes on the same chain as atomic arbitrage, since these arbitrages can be carried out within the same transaction, whereas non-atomic arbitrage stems from price discrepancies involving DEXes on two different chains or, presumably more common, centralized exchanges. The prior is known as cyclic arbitrage and has been extensively studied [1]. We focus on the latter, which is naturally more challenging to analyze as it involves actions outside the respective blockchain. We note that while non-atomic arbitrage opportunities exist on various chains we focus on Ethereum, as it is home to the by far largest portion of DeFi applications [2].
@@
In this work, we present the first in-depth analysis of non-atomic arbitrage on Ethereum DEXes to show that non-atomic arbitrage, a type of MEV, accounts for a significant proportion of the total DEX volume. We further identify alarming centralizing trends in the block building market as a direct consequence of non-atomic arbitrage.
@@ 2.3 Decentralized Exchanges
DEXes are one of the most popular DeFi applications. They allow users to exchange cryptocurrencies on the blockchain without giving up custody of their assets. While traditional CEXes generally utilize a limit order book to match orders, most DEXes are constant function market makers (CFMM). With a CFMM, users execute their order against a liquidity pool holding reserves of two or more cryptocurrencies. Trades against the pool execute according to a pre-defined trading function specified in the pool’s smart contract, i.e., the pool only accepts trades that keep its trading function constant. Importantly, every trade that executes in a CFMM pool moves the pool’s price deterministically, dependent only on the pool’s reserves ahead of the trade and the trade’s input size. Consequently, if an arbitrage opportunity arises, the assets locked in the liquidity pool determine how large the arbitrage trade can be. The larger the liquidity pools, the larger the required arbitrage trades are to synchronize the prices across different trading venues.
@@ 3.1 Ethereum Blockchain
Maximal Extractable Value. To differentiate non-atomic arbitrage trades from types of MEV and to compare their respective prevalence on DEXes, we collect MEV transaction labels from zeromev [9]. These labels specify whether a transaction was part of a sandwich attack (i.e., victim and attacker transaction), performs a cyclic arbitrage, or a liquidation.
@@ 3.2 PBS Relays
We collect data from eleven relays: Aestus, Agnostic, Blocknative, bloXroute (Ethical), bloXroute (Max Profit), bloXroute (Regulated), Eden, Flashbots, Manifold, Relayoor and UltraSound. The relays implement public APIs that give access to the blocks delivered by the relays to the proposers, as well as the bids received by the builders. Note that some relay APIs, namely Blocknative, bloXroute (Ethical), Eden, and Relayoor, were not reachable. For these, our data only extends to 31 April 2023, i.e., the last time we scraped the data. Further, bloXroute (Ethical), bloXroute (Max Profit), and bloXroute (Regulated) removed access for builder bids from their APIs. When querying the builder bids API for blocks that passed through the respective bloXroute relay, the response is consistently empty. Thus, for the three bloXroute relays the bid data only extends to 31 January 2023, i.e., the last time we scraped that data and received non-empty responses.
@@ 3.4 Cryptocurrency Price
4 Non-Atomic Arbitrage on DEXes Model
We illustrate the process a non-atomic arbitrageur follows to get the on-chain leg of the trade included in Figure 2. Upon detecting an arbitrage opportunity, the searcher submits the desired transaction to the builder. The builder then builds their block and sends it together with their bid to the relay. Next, if requested, the relay sends the highest value block to the proposer who in turn picks the highest bidding block to propose to the network. Importantly, even though visualized separately one entity could for instance operate a builder and a searcher. In this case, we would refer to the latter as an integrated searcher.
@@ 4.1 Non-Atomic Arbitrageur Model
In this section we develop a model for the profit of a non-atomic arbitrageur between two currencies $X$ and $Y$. As discussed in Section 2.3, most DEXes are a CFMM with the largest DEX being Uniswap [8]. Uniswap currently has two active versions, V2 [14] and V3 [15]. Here, we focus on the V2 version, which is also employed by the DEX Sushiswap [16], but the model also holds for V3 between two ticks. The Uniswap CFMM imposes, $x\cdot y=L^{2}$, where $x$ and $y$ stand for the number of $X$- and $Y$-tokens locked in the liquidity pool, and $L$ is referred to as the liquidity. The (marginal) price $P$ for $X$-tokens in terms of $Y$-tokens is given as $P=y/x$. The reserves can be expressed in terms of the liquidity and price by solving for $x$ and $y$, yielding $x=L/\sqrt{P}$, $y=L\sqrt{P}$.
The price on a DEX on Ethereum is $P_{\mathrm{on}}$ and the average off-chain price is $\tilde{P}_{\mathrm{off}}$, i.e., the average price received for an order. [^1] The average price $\tilde{P}_{\mathrm{off}}$ can contain transactions on multiple off-chain exchanges. In general, it is the best price attainable for a transaction of a given size. We denote the price difference between the off-chain and on-chain as $\Delta P=\tilde{P}_{\mathrm{off}}-P_{\mathrm{on}}$. Without loss of generality, we assume the searcher discovers $\Delta P>0$, i.e., the price of $X$-tokens off-chain is higher such that the arbitrageur buys $X$-tokens on the DEX. During the trade, they pay $\Delta y$ of which $f\cdot\Delta y$ is paid in fees to the liquidity providers, where $f$ is the protocol fee. With this and the constant product $L^{2}=xy$, the amount of $X$-tokens the trader receives is given by
@@
Apart from $f$, the arbitrageur is charged a fee by the off-chain exchanges which we denote as $g$.
As long as the off-chain price remains higher than the on-chain price, the arbitrageur can extract more value by buying more $\Delta x$ on the DEX. Thus, the arbitrageur will trade on the DEX until the price they receive for an infinitesimal unit of $X$ is $\tilde{P}_{\mathrm{off}}(1-g)$. [^2] Here we assume that the off-chain price is unaffected by the arbitrage. In practice, the arbitrageur would receive a quote for a given trade size and can be used as $\tilde{P}_{\mathrm{off}}$. The pool’s marginal price at the end of the trade is then given by
$P_{\mathrm{end}}=\tilde{P}_{\mathrm{off}}(1-g)(1-f)=\tfrac{y+(1-f)\Delta y}{x-\Delta x}.$
Simultaneously, to this buy order, the arbitrageur sells the same amount of $X$-tokens, $\Delta x$ through the off-chain exchanges at an average price of $\tilde{P}_{\mathrm{off}}$.
Thus, the arbitrageur receives
$\Delta y_{\mathrm{off}}=\tilde{P}_{\mathrm{off}}(1-g)\Delta x=\tfrac{x\Delta y}{x-\Delta x},$
@@
$\Delta y_{\mathrm{off}}-\Delta y=L\tfrac{\left(\sqrt{P_{\mathrm{on}}}-\sqrt{(1-f)(1-g)(\Delta P+P_{\mathrm{on}})}\right)^{2}}{(1-f)\sqrt{P_{\mathrm{on}}}}$
where we also expressed $x$ and $y$ in terms of $L$ and the starting on-chain price $P_{\mathrm{on}}$. The profit as a function of the price change $\Delta P$ is plotted in Figure 3. The profit becomes significant for larger price changes, showing that sufficient price difference will attract non-atomic arbitrage trades. Note that profitable arbitrage only exists for $P_{\mathrm{on}}<\tilde{P}_{\mathrm{off}}(1-f)(1-g)$, as otherwise the fees paid to the liquidity pool exceed the price difference that can be profited from.
Figure 3: Arbitrageur profit as a function of the price difference for $L=10^{6}$, $P_{\mathrm{on}}=1$, $f=0.3\%$, and $g=0.1\%$. Note we only plot the profit in the profitable range.
We further note that while an arbitrageur still needs to pay the builder to have their trade included, the result shows how valuable the arbitrage trade is for the searcher and limits the bid they would be willing to submit.
@@ 4.2 Case study of Block 18,360,789
To provide a better understanding of non-atomic arbitrage, we go through a case study of block 18,360,789 – a block with a significant price change in the lead-up to the block, i.e., the time between the previous block proposal and the block proposal itself. In Figure 4, we plot the time and the value of bids from the builders. Note that we hightlight bids from rsyncbuilder (shown in red) and beaverbuild (shown in blue), as these two builders were previously identified as having integrated searchers that perform these non-atomic arbitrage trades [4]. The bids from all remaining builders are shown in yellow. We further indicate the time of the bid corresponding to the block that was chosen by the proposer, i.e., the relays deliver the highest bid to the proposer when the proposer requests it. Finally, we also plot the relative price change of ETH-USDT and BTC-USDT on Binance.com during the same time. Recall that USDT is a stablecoin pegged to the \$.
One immediately notices that the time in the lead-up to the block is characterized by extremely high price volatility of both ETH and BTC in comparison to USDT. The relative price change exceeds 0.6% and 1.3% in less than ten seconds respectively. This price change creates an arbitrage opportunity, as the prices on DEXes do not move in between blocks. Thus, as outlined in the previous section, the first transaction in a DEX pool can close this price difference and thereby profit from the arbitrage opportunity.
@@
tx index searcher pool tokenIn tokenOut amount [\$] 0 rsyncsearcher3 0x88e6a...f5640 USDC ETH 2148195.308 1 rsyncsearcher3 0x99ac8...abc35 USDC BTC 659507.701 2 rsyncsearcher3 0x4585f...a20c0 ETH BTC 554524.411 3 rsyncsearcher3 0x9db9e...8425b USDT BTC 186858.044 4 rsyncsearcher3 0x11b81...697f6 USDT ETH 523225.597 5 rsyncsearcher3 0x4e68c...dfa36 USDT ETH 524150.487 6 rsyncsearcher3 0xcbcdf...d62ed ETH BTC 495617.185 8 searcher1 0x8ad59...6e6d8 USDC ETH 562787.652 9 rsyncsearcher3 0x60594...5a270 DAI ETH 116090.952 10 rsyncsearcher3 0xc2e9f...a25f8 DAI ETH 105871.055 12 searcher1 0xd51a4...aae46 USDT BTC 17463.007 14 rsyncsearcher3 0x391e8...86649 DAI BTC 9214.814 17 rsyncsearcher3 0x06da0...84553 USDT ETH 18283.993 18 rsyncsearcher3 0xc3d03...5882f DAI ETH 10533.977 19 rsyncsearcher3 0x397ff...7aca0 USDC ETH 14457.825 21 rsyncsearcher3 0xceff5...d3a58 ETH BTC 15450.829 22 rsyncsearcher3 0x824a3...c0673 ETH BLUR 11451.829 23 rsyncsearcher3 0x56534...a83b2 USDT BTC 4822.975 27 rsyncsearcher3 0xcbfb0...7ef3c USDC BTC 1538.249 28 rsyncsearcher3 0x290a6...f5b42 ETH MATIC 14629.722 31 searcher5 0xcd828...a9fbf FXS ETH 6247.832 33 rsyncsearcher3 0x919fa...daf79 ETH CRV 4954.943 35 rsyncsearcher3 0x34704...08d60 USDT UNI 1106.016 36 rsyncsearcher3 0xa3f55...dfd74 LDO ETH 9727.578 43 rsyncsearcher3 0x9febc...c82ae USDC 1INCH 516.780 44 rsyncsearcher3 0x1d420...bd801 ETH UNI 6336.821
@@
TABLE I: Non-atomic arbitrages identified in block 18,360,789 on 16 October 2023. There are 26 non-atomic arbitrage swaps with a total volume of \$6,023,565 executed by merely three searchers (cf. [17] for a mapping from address to searcher name). Further, rsyncbuilder’s integrated searcher is responsible for 23 of these arbitrages. We indicate the tx index, the searcher, pool, and the tokenIn (the token that was sold), the tokenOut (the token that was bought), and the amount (the trade volume in \$).
@@
Coming back to Figure 4, we can observe that around five seconds after the price starts to change on Binance.com the bids start to increase. At this point, the price difference appears to be big enough for non-atomic arbitrage to be profitable. Further, we find that bids from the builders that are associated with non-atomic arbitrage transactions are higher than those from the rest and that the bids continue to increase as the ETH and BTC prices on Binance.com increase. The block chosen by the proposer was built by rsyncbuilder and its bid, i.e., the value received by the proposer was a staggering 10.32 ETH ($\approx$\$23,000 at the time of this writing). In Table I, we take an in-depth look a the non-atomic arbitrage trades we identified with our heuristics that we will introduce in the following Section 5. In the block, there were 26 transactions with volume exceeding \$6 million that we identified as performing non-atomic arbitrage trades. Remarkably, all but three of these transactions were by the rsyncsearcher3 – a searcher we identified to be linked to the rsyncbuilder that won the block (cf. Section 6.3). Additionally, we highlight that the rsyncsearchers paid a remarkable 11.15 ETH (i.e., more than the PBS bid) for their transactions. Thus, the rsyncbuilder built a high-value block largely attributed to fees paid by trades we labeled as non-atomic arbitrage by its own integrated searcher.
@@
Note that most of trades buy ETH or BTC for a stablecoin. As we saw in Figure 4 the prices of these two cryptocurrencies rose on Binance.com, and they were thus trading for cheaper on DEXes at the beginning of the block. For example, if we look at the price at the beginning of the block, we see that the price of the ETH-USDT Uniswap V3 pool was 1563.57, and at the end of the block, it was 1574.15. The Binance.com price was 1564.61 at the beginning of the block’s slot, while it rose to 1574.63 by the time the block was proposed. Notice that the arbitrageurs drove the price almost exactly to the off-chain price, as we show to be optimal in our previous analysis (cf. Section 4.1). Additionally, some transactions exchange cryptocurrencies for cryptocurrencies (i.e., no stablecoins), such as the transaction at index 2 and 6, which exchange ETH for BTC. The relative price increase measured in USDT of BTC was higher than that of ETH. Thus, BTC relative to ETH was also available for cheaper on DEXes at the beginning of the block. Hence, the searcher also swaps ETH for BTC.